You observe the following sample: {2.3, 4.7, 2.9, 4.0, 3.2, 7.8, 9.1, 3.2, 2.4}. Calculate the sample mean and the sample variance.
Answer #1

sample mean = \sum Xi / n

= (2.3+ 4.7+ 2.9 + 4.0+ 3.2+ 7.8+ 9.1+ 3.2+ 2.4) / 9

= 39.6 / 9

= 4.4

Sample mean = 4.4

Sample variance = \sum Xi2 - n * (sample mean)2 / n-1

= (2.32+ 4.72+ 2.92 + 4.02+ 3.22+ 7.82+ 9.12+ 3.22+ 2.42) - 9 * 4.42 / 8

= 5.93

Sample variance = 5.93